HUBBLE SPACE TELESCOPE DAILY REPORT #1992 PERIOD COVERED: 0000Z (UTC) 10/27/97 - 0000Z (UTC) 10/28/97 Daily Status Report as of 301/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed WF/PC-2/NIC/3 7382 (The Star-Formation History of the Large Magellanic Cloud) The WF/PC-2 and the Near Infrared Camera (Camera Three) were used to make observations of the disk of the Large Magellanic Cloud. These data will enable us to quantitativly study the intermediate to low-mass stellar populations and the star-formation history of the LMC. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed STIS/CCD 7782 (STIS Non-Scripted Parallel Archive Proposal - Half Orbit) The Space Telescope Imaging Spectrograph (CCD) was used for a non-scripted parallel archive proposal (one-half orbit). The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed Three Sets of WF/PC-2 7621 (WF/PC-2 Cycle 7 Supplemental Darks PT1/3) The WF/PC-2 was used to make supplemental dark calibration images. This set of calibrations is intended to obtain three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.4 Completed STIS/CCD 7802 (Daily Darks to Update Acquisition Bad Pixel Table) The Space Telescope Imaging Spectrograph (CCD) was used to make calibration observations to obtain daily darks to update the acquisition bad pixel table. These daily darks are intended to identify hot pixels for update to the acquisition on-board FSW hot pixel table. Daily hot pixel lists can also be used for science data analysis. The observations were executed as scheduled, and no anomalies were noted. 1.5 Completed STIS/CCD 7634 (CCD Flat-Field Stability and Cosmetic Defect Fraction) The Space Telescope Imaging Spectrograph (CCD) was used to investigate and monitor the STIS CCD imaging flat-field stability and cosmetic defects over monthly periods though out Cycle 7. A bias calibration observations was also taken. The observations were executed as scheduled, and no anomalies were noted. 1.6 Completed Six Sets of WF/PC-2 7620 (WF/PC-2 Cycle 7 Standard Darks) The WF/PC-2 was used to make a series of Cycle 7 standard dark calibration images. This dark calibration program obtains dark frames every week in order to provide data for the ongoing calibration of the CCD dark current rate, and to monitor and characterize the evolution of hot pixels. . The third iteration of this proposal was affected by the guide star acquisition failure at 300/0634Z (and related failures) reported in HSTAR 6221 and in yesterday's report. Otherwise, the remaining observations were executed as scheduled, and no further anomalies were noted. 1.7 Completed Four Sets of STIS/MA2 7604 (Cycle 7 MAMA Dark Measurements) The Space Telescope Imaging Spectrograph (MA2) was used to make MAMA dark measurements. These darks will be used in later image reductions and also trended over the instrument's life time as an indicator of the detector's degradation or damage. The third iteration of this proposal was not executed since STIS was in a suspend state. Otherwise, the observations were completed as scheduled, and no further problems were reported. 1.8 Completed NIC/1/2 7180 (Surface and Atmospheric Features on Titan) The Near Infrared Camera (Cameras One and Two) was used to make high resolution infrared observations of the surface and atmospheric features on Saturn's moon Titan. The primary objective is to measure the wavelength dependence of discrete surface features to investigate possible compositional heterogeneity of Titan's surface. The observations were executed as scheduled, and no anomalies were noted. 1.9 Completed FGS 6878 (Parallax of NGC-6853) The Fine Guidance Sensors were to be used used to make astrometric observations of stars located in NGC-6853. This proposal was affected by the guide star acquisition failure at 300/0634Z (and related failures) reported in HSTAR 6221 and in yesterday's report. 1.10 Completed NIC/1/2/3 7726 (NICMOS One Orbit Parallel) The Near Infrared Camera (Cameras One, Two and Three) was to be used to make a one orbit parallel calibration observations. . This proposal was affected by the guide star acquisition failure at 300/0634Z (and related failures) reported in HSTAR 6221 and in yesterday's report. 1.11 Completed WF/PC-2/NIC/3/STIS/CCD 7563 (Francis Cluster) The WF/PC-2 and the Near Infrared Camera (Camera Three) was to be used to make observations of the cluster Gal-Clus-2 (a.k.a., Francis Cluster) using various modes. This proposal was affected by the guide star acquisition failure at 300/0634Z (and related failures) reported in HSTAR 6221 and in yesterday's report. This proposal was also affected by STIS being in a suspend state. 1.12 Completed NIC/2 7420 (A Near-Infrared Search for Very Low Mass Companions to Stars Within 10 Parsecs of the Sun) The Near Infrared Camera (Camera Two) was used to make searching observations for very low mass companions to variable star HD157881. This star is known to be within 10 parsecs of the Sun. The observations were executed as scheduled, and no anomalies were noted. 1.13 Completed FOC/96 7685 (FOC F/96 UV Throughput Monitoring) The Faint Object Camera (f/96) was used to make observations of standard star BPM16274 in order to perform ultraviolet throughput monitoring. These observations are taken four time during the cycle. An internal calibration flat image was also taken. The observations were completed as scheduled, and no problems were reported. 1.14 Completed STIS/CCD 7783 (STIS Parallel Science) The Space Telescope Imaging Spectrograph (CCD) was to be used to make observations taken in parallel with the prime science. This proposal was not executed since STIS was in a suspend state. 1.15 Completed Three Sets of WF/PC-2 7786 (Pure Parallel Archive Survey - Continuation) The WF/PC-2 was used to make observations as part of an archival survey program in which the images were taken in parallel with the prime science observations. These WF/PC-2 parallel observations were obtained in the ultraviolet in a search for interesting objects. The observations were completed as planned, and no problems were reported. 1.16 Completed NIC/2 7229 (High Redshift Gravitational Lens Galaxies) The Near Infrared Camera (Camera Two) was used to make observations of the far infrared shifted gravitationally lensed galaxy FSC10214. Its high interstellar medium mass and high far-infrared luminosities suggest that it may also have massive protogalaxies undergoing initial galactic starbursts. A dark calibration image was also taken. No anomalies were noted. 1.17 Completed NIC/1/2/3 7421 (The Relation Between Quasar Luminosity and Host Galaxy Mass) The Near Infrared Camera (Cameras One, Two, and Three) was used to obtain ground-based, near-infrared imaging of nearby quasar PG1121+422 which has indicated a relation between the luminosity of the nucleus and the luminous mass of the host galaxy. The observations completed as planned, with no reported problems. 1.18 Completed NIC/1/2 7320 (Distribution of Forbidden Iron [Fe-II] Emission at the Galactic Center) The Near Infrared Camera (Cameras One and Two) was used to make observations of the distribution of the forbidden Iron (Fe-II) emissions from the center of our Milky Way galaxy. A conspicuous feature known as the mini-cavity appears as a hole in the distribution of ionized gas within a few arcseconds of the dynamical center of the Galaxy. The observations were executed as scheduled, and no anomalies were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 9 Successful : 9 Scheduled Reacquisitions : 7 Successful : 7 2.2 FHST Updates: Scheduled : 18 Successful : 18 2.3 Operations Notes: The Battery 1 Capacity test was completed with the return to the six battery configuration at 300/1853Z. STIS was successfully recovered from suspend at 300/1853Z. The CCD was set to operate. While in the suspend state, proposals described in 1.7, 1.11, and 1.14 were affected. A NSSC-1 patch was performed at 300/2241Z, and both MAMAs were commanded to low voltage at that time. STIS operations resumed normally. The STIS Bad Pixel table was uplinked at 300/2311Z. Memory monitors were set up for STIS MCE operations at 300/2321Z. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH